OpenStudy (anonymous):

how many photons with 10 ev are required to produce 20 joules of energy A. 2000 B. 3.2x10^18 C. 1.2X10^19 D. 3.3x10^19 E. 1.2X10^20

OpenStudy (anonymous):

1eV=1.6*10^-19J no of photons = total energy/energy in 1 photon

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

well, the energy of 1 photon is =10eV =10*1.6*10^-19 J =1.6 *10^-18J And total energy to be produced is given = 20J So how many photons would u require to produce this much energy. Just divide the total energy to be produced by energy associated with a single photon.